OpenStudy (anonymous):

Factor: 9x^2-468x+1944

OpenStudy (anonymous):

easier to look at: 9(x^2 -52 +216)

OpenStudy (anonymous):

yeah that's what i did, still.... plz help!

OpenStudy (anonymous):

and thx for replying

OpenStudy (anonymous):

actually it can't be factored

OpenStudy (anonymous):

yeah thats what i thought, but I need to solve for x.... so idk what to do

OpenStudy (anonymous):

you can use the quadratic formula on your second order polynomial of the form ax^2 + bx + c \[x_{1and2} = \frac{ -b \pm \sqrt{b^2 - 4ac} }{ 2a }\]

OpenStudy (anonymous):

ok thanks!